# ASDF: Assembly State Detection Utilizing Late Fusion by Integrating 6D Pose Estimation

### v2

@@ 2.1.3 Pose2State Module
Given the inherent co-dependency between pose and state, we harness this relationship within our Pose2State module. This module is designed to forecast the probability of the assembly state at time . Our Pose2State module seamlessly integrates two key components: the deep learning-based prediction and the prediction derived from the pose-based assembly . By combining these predictions, our module effectively determines the final assembly state, leveraging both the insights from deep learning techniques and the pose-based assembly analysis.
For the final assembly state we consider the time component and leverage the previously stored assembly state . The combination of the individual predictions is then weighted:
This results in the final state calculation:
@@ 2.2 Assembly State Dataset
To train ASDF, we created a synthetic dataset using 3D printable parts to facilitate real-world testing. We utilized the 3D printable parts from the GBOT dataset [19] as a foundation. In addition to that existing reproducible setup [19], we introduced assembly state information and incorporated hand occlusion during training to enhance the model’s robustness. Our dataset adheres to the camera specifications of the Azure Kinect DK, with a resolution set to pixels. The camera is positioned in a top-down view to simulate a capturing setup commonly found in medical or industrial scenarios.
@@ 2.2.1 Synthetic Dataset
Training Dataset
Our synthetic dataset features 6D pose estimation and assembly state detection ground truth data. It is generated using BlenderProc [7] and incorporates several domain randomization techniques, including distractors, randomized noise, lighting variations, occlusions, and background changes [1, 29]. Domain randomization plays a vital role in synthetic data generation [29, 1]. Therefore, we introduce distracting objects [12, 17] and simulate hand occlusion to mimic real-world scenarios, as illustrated in Figure 2 and Figure 3.
@@
We adhere to the defined assembly process for each assembly object by generating a continuous set of assembly images for each object. We maintain the presence of distracting objects and incorporate hands at realistic assembly positions. The objects are randomly distributed within the camera’s field of view. We consistently introduce hand occlusion in the evaluation set and provide pose and state information for each image.
A key aspect of our test set is the inclusion of incorrectly assembled states. In a real-world assembly scenario mounting together parts with missing pieces in between can happen especially if the parts are not necessarily connecting parts. Therefore, we added incorrect assembly states. Since each object contains various assembly states, the number of evaluation images per assembly object varies, as illustrated in Table 1. This variability ensures comprehensive testing of the model’s ability to detect and handle different assembly scenarios.
@@ 3.2 Metrics
We evaluate our system based on two primary aspects: 6D pose estimation and state detection. For the pose prediction we report the absolute translation (), rotation error and Average Distance Error (ADD)(S) [11]. ADD(S) describes the average distance error for unsymmetric (ADD) and symmetric (ADD(S)):
is a vertex point from the 3D object mesh, represent the number of vertices of the object mesh.
Following the definition of Hodan et al. [13], we calculate the translation error according to:
Where represents the error between the ground truth translation and the predicted translation .
The rotation error () is calculated according to the following formula:
denotes the ground truth rotation and the ground truth translation. describes the predicted rotation and the predicted rotation.
For the sate detection we calculate the F1 score. The F1 score consist of calculation Precision and Recall for each state () and each assembly part ().
Since we target AR to utilize our state and pose detection, we additionally evaluate the runtime. For the runtime we calculated the mean over all assembly objects.
@@
3.3 Datasets
| Method | Network | Runtime [ms] | etrans [mm] | erot [°] |
| No refinement | ||||
| YOLOv8Pose + state | n | 24.83 | 29.13 | 18.76 |
| YOLOv8Pose + state | xl-p6 | 50.83 | 19.11 | 11.14 |
| Translation refinement | ||||
| ASDF-mini | n | 32.45 | 8.13 | 18.76 |
| ASDF + Seg | n | 64.46 | 8.52 | 18.76 |
| ASDF + Seg | xl-p6 | 85.03 | 6.53 | 11.14 |
| ICP refinment | ||||
| ASDF + ICP | n | 151.77 | 12.27 | 18.96 |
| ASDF +ICP | xl-p6 | 145.76 | 9.24 | 14.35 |
| Ours | ||||
| ASDF | xl-p6 | 55.70 | 5.98 | 11.14 |
| Network | Runtime [ms] | etrans [mm] | erot [°] |
| n | 24.83 | 29.13 | 18.76 |
| m | 33.72 | 27.85 | 15.21 |
| l | 35.55 | 23.69 | 13.09 |
| xl-p6 | 50.83 | 19.11 | 11.14 |
GBOT Dataset:
The GBOT dataset [19] is a exclusively a synthetic dataset designed for 6D object pose/tracking tasks. It comprises 3D printable assembly parts and contains five evaluation sequences, each presenting different levels of difficulty for object tracking (normal, dynamic, hand occlusion, and blur). Utilizing the GBOT dataset, we conduct a comprehensive comparison of our approach against state-of-the-art methods in 6D pose estimation performance, particularly focusing on assemblies with dynamically changing states.
ASDF Dataset:
The ASDF dataset provides ground truth annotations for 6D object poses and assembly states of 3D printing parts. It includes four distinct assembly parts and varying numbers of images as outlined in Table 1. Within the context of the ASDF dataset, we assess the performance trade-offs, evaluate the effectiveness of our pose refinement module, and analyze the accuracy of our assembly state detection approach.
3.4 ASDF Results
3.4.1 Performance Tradeoff
Since ASDF leverages object pose information to refine the state, our initial focus lies on evaluating the pose performance, specifically concentrating on rotation and translation errors. As depicted in Table 4, it is evident that the fastest network performance can be attained using the smallest network configuration, which is a rational expectation. Notably, networks m and l exhibit similar performance, whereas the xl-p6 architecture, serving as the underlying framework, demonstrates the most favorable results in terms of rotation and translation errors.
@@ 3.4.2 Pose Refinement
As n shows the best Runtime and xl-p6 the best performance in Table 4, we compare these two networks using different refinement methods (refinement w/o seg, w seg and ICP).
The improvement of our translation refinement compared to the pure network-based output can be seen in Figure 6. In terms of numbers, the result can be seen in in Table 4.
ASDF builds upon the use of keypoints for the 6D pose estimation. However, we, considered other refinement methods as well to lift the pose performance. Since semantic segmentation provides clearer object boundaries, we compared our keypoint-based approach using a refinement step with a second YOLOv8 semantic segmentation network.
3.4.3 Assembly State Detection
As shown in Table 4 and Table 3, ASDF shows promising results in terms of 6D pose estimation. However, the main focus of ASDF is the 6D pose estimation and assembly state detection. Therefore, we compare pure deep learning-based 6D pose estimation and ASDF on the ASDF dataset.
Comparing the assembly state detection using the F1 score, see Table 2, shows that ASDF can predict the assembly state better (79.52) compared to a pure deep learning-based approach (73.85). Moreover, as denoted in Table 2, the 6D pose performance shows an improved performance by outperforming the deep learning-based approach in ADD(S) (95.74 (ours) compare to 84.92). Moreover, for the translation error we can reduce the average error by 14.2 mm.
@@ 3.4.3 Assembly State Detection
| Asset | Condition | |||||||||||||||||||||
| Normal | 100.0 | 1.8 | 3.0 | 91.5 | 9.3 | 3.8 | 89.8 | 4.7 | 25.8 | 100.0 | 0.1 | 38.7 | 100.0 | 1.3 | 38.7 | 100.0 | 0.6 | 2.1 | 99.5 | 0.7 | 2.7 | |
| Corner | Dynamic | 100.0 | 1.8 | 2.7 | 99.0 | 2.5 | 4.8 | 88.4 | 3.7 | 25.0 | 100.0 | 1.5 | 46.8 | 100.0 | 2.0 | 47.8 | 100.0 | 0.6 | 23.2 | 100.0 | 0.5 | 3.5 |
| Clamp | Hand | 68.6 | 8.3 | 38.4 | 45.4 | 54.1 | 97.1 | 66.6 | 8.8 | 37.1 | 68.4 | 7.4 | 59.4 | 68.4 | 7.5 | 58.3 | 81.9 | 5.7 | 90.4 | 90.6 | 4.4 | 84.7 |
| Blur | 100.0 | 1.9 | 3.4 | 97.3 | 4.0 | 4.3 | 88.9 | 5.0 | 30.4 | 100.0 | 0.5 | 2.2 | 100.0 | 1.0 | 3.0 | 100.0 | 0.6 | 2.1 | 99.9 | 0.6 | 2.1 | |
| Normal | 100.0 | 0.4 | 4.7 | 99.6 | 1.4 | 10.1 | 90.6 | 1.8 | 5.4 | 100.0 | 0.2 | 2.6 | 100.0 | 0.3 | 3.0 | 100.0 | 0.2 | 2.1 | 100.0 | 0.5 | 3.3 | |
| Geared | Dynamic | 100.0 | 0.3 | 4.1 | 99.9 | 1.3 | 9.3 | 92.7 | 1.4 | 9.9 | 100.0 | 0.2 | 2.5 | 100.0 | 0.4 | 3.6 | 100.0 | 0.2 | 2.3 | 100.0 | 0.5 | 3.6 |
| Caliper | Hand | 100.0 | 0.8 | 6.9 | 99.2 | 2.4 | 12.5 | 96.5 | 0.9 | 4.4 | 85.4 | 3.0 | 30.0 | 85.5 | 3.1 | 30.0 | 85.4 | 3.0 | 30.0 | 99.6 | 0.8 | 7.0 |
| Blur | 100.0 | 0.4 | 4.9 | 99.6 | 1.4 | 9.9 | 98.9 | 0.9 | 8.7 | 100.0 | 0.2 | 2.5 | 100.0 | 0.3 | 2.8 | 100.0 | 0.2 | 2.2 | 100.0 | 0.5 | 3.6 | |
| Normal | 100.0 | 1.4 | 4.8 | 71.1 | 8.8 | 4.8 | 74.1 | 6.6 | 13.8 | 89.8 | 2.1 | 16.7 | 89.4 | 2.3 | 16.5 | 99.8 | 0.6 | 7.2 | 93.8 | 2.4 | 3.6 | |
| Nano | Dynamic | 90.0 | 1.6 | 5.0 | 72.6 | 7.6 | 4.6 | 63.4 | 10.3 | 15.6 | 87.3 | 3.9 | 15.3 | 87.3 | 4.1 | 15.8 | 96.0 | 2.4 | 20.0 | 92.9 | 2.5 | 3.7 |
| Vise | Hand | 98.6 | 1.9 | 5.3 | 65.9 | 11.0 | 4.7 | 61.4 | 13.6 | 15.3 | 76.5 | 6.0 | 18.3 | 75.8 | 6.1 | 18.1 | 72.9 | 8.7 | 14.5 | 87.8 | 3.1 | 7.3 |
| Blur | 100.0 | 1.5 | 5.4 | 70.9 | 9.9 | 5.2 | 61.9 | 11.6 | 15.2 | 91.6 | 1.9 | 11.4 | 91.5 | 2.1 | 11.1 | 95.7 | 0.7 | 25.1 | 92.7 | 3.0 | 4.7 | |
| Normal | 93.5 | 1.1 | 8.6 | 73.5 | 7.7 | 17.6 | 86.5 | 4.7 | 8.9 | 96.0 | 1.2 | 1.1 | 95.9 | 1.4 | 2.3 | 98.8 | 0.4 | 0.9 | 83.7 | 3.0 | 4.7 | |
| Screw | Dynamic | 94.5 | 1.1 | 9.4 | 82.0 | 6.7 | 18.4 | 86.4 | 5.6 | 27.0 | 95.9 | 1.7 | 2.1 | 95.9 | 2.2 | 3.4 | 98.8 | 0.6 | 1.7 | 91.6 | 2.1 | 6.1 |
| Clamp | Hand | 92.1 | 1.8 | 13.5 | 67.2 | 14.2 | 39.8 | 60.1 | 14.3 | 37.7 | 73.4 | 6.9 | 53.4 | 73.1 | 7.1 | 53.4 | 68.7 | 7.9 | 61.9 | 83.9 | 4.9 | 27.2 |
| Blur | 96.1 | 0.9 | 8.7 | 85.6 | 6.5 | 18.6 | 86.5 | 5.6 | 34.6 | 95.7 | 3.0 | 6.6 | 95.7 | 3.2 | 7.7 | 98.6 | 1.2 | 1.1 | 91.1 | 3.0 | 4.8 | |
| Mean | Overall | 95.8 | 1.7 | 8.1 | 82.52 | 9.3 | 16.59 | 80.8 | 6.2 | 19.7 | 91.3 | 2.5 | 19.4 | 91.2 | 2.8 | 19.7 | 93.5 | 2.1 | 17.9 | 94.2 | 2.0 | 10.8 |
@@ 3.5 GBOT Results
Since our system addresses 6D object pose estimation and assembly-state detection existing approaches face the limitation of being either used for state detection combined with 2D object detection [31] or are limited to non accessible datasets [24, 37]. However, in terms of 6D pose estimation and 6D object tracking some approaches can handle 6D pose estimation for assembled parts.
Given that we utilize the same objects as Li et al. [19], we assess our trained models from the ASDF dataset on the GBOT dataset without the need for retraining on their training set. As illustrated in Table 5, our approach demonstrates superior performance compared to GBOT and their fusion of GBOT with deep learning-based re-initialization of object tracking. Moreover, in comparison to state-of-the-art tracking methods, Li et al. [19] introduce a deep learning-based approach for 6D pose estimation, which shares the same base architecture as our method. The outcomes presented in Table 5 shows that our approach, with integrated state detection during both training and inference, achieves better results than a pure deep learning-based approach.

### v3

@@ 2.1.3 Pose2State Module
Given the inherent co-dependency between pose and state, we harness this relationship in our Pose2State module in a late fusion manner. This module is designed to forecast the state probability of the assembly state at time . Our Pose2State module seamlessly integrates two key components: the deep learning-based prediction and the prediction derived from the pose-based assembly . By combining these predictions, our module effectively determines the final assembly state, leveraging both deep learning and the pose-based assembly state.
We first combine the deep learning-based state prediction and the pose-based state prediction :
For the final assembly state we consider the time component and leverage the previously stored assembly state (). represents the average normalised probability of the two neighbouring states from the previous time step .
This results in the probability () for the final assembly state at time . We fuse the state probability () our deep learning-based and relative-pose based estimation () and weight them:
@@ 2.2 Assembly State Dataset
To address assembly states and 6D pose estimation, we created a synthetic dataset222GitHub ASDF with 3D-printed parts to facilitate testing under real conditions. We used the 3D printable parts from the GBOT dataset [21] as a foundation. In addition to the existing reproducible setup [21], we introduced assembly state information, wrong assembly states and incorporated hand occlusion during training to enhance the model’s robustness. Our dataset adheres to the camera specifications of the Azure Kinect DK, with a resolution set to pixels. The camera is positioned in a top-down view to simulate a capturing setup commonly found in medical or industrial scenarios.
@@ 2.2.1 Synthetic Dataset
Training and Validation Dataset
Our synthetic dataset includes 6D pose estimation and assembly state detection ground truth data with 20k images per assembly, using an 80:20 training split. For testing each test scene contains a various number of images with several levels of domain randomization, see Table 1. Our dataset is generated using BlenderProc [6]. Domain randomization is crucial in synthetic data generation [29, 1]. Thus, we introduce distracting objects [13, 18] and simulate hand occlusion to reflect real-world scenarios, as shown in Figure 3. In addition to, to distracting objects and hand occlusion, we add randomized noise, different lighting variations, and varying background materials [1, 29].
@@
For each assembly we generate a continuous sequence of assembly images resulting in a full assembly video. The test scenes contain one full assembly sequence including error states to test the robustness of the assembly state detection. We maintain the presence of distracting objects and incorporate hands at realistic assembly positions to mimic occlusion and test robustness. The objects are randomly distributed within the camera’s field of view. Each frame contains assembly state and 6D pose ground truth data.
A key aspect of our test set is the inclusion of incorrectly assembled states. In a real-world assembly scenario mounting together parts with missing pieces in between can happen especially if the parts are not necessarily connecting parts. Therefore, we added incorrect assembly states to test robustness. Since each object contains various assembly states, the number of evaluation images per assembly object varies, as illustrated in Table 1. This variability ensures comprehensive testing of the model’s ability to detect and handle different assembly scenarios.
@@ 3 Evaluation
To benchmark the accuracy of the assembly state detection and 6D pose estimation, we evaluate our approach on our ASDF dataset and on the GBOT [21] datasets. Other comparisons with existing assembly state detection approaches [32, 23] in 2D are not possible, since both pose and state are relevant for the final prediction. Other works focus purely on state changes [31], purely on 6D pose performance [21] or build upon non publicly available [38, 25] assets or data.
GBOT Dataset:
The GBOT dataset [21] is a exclusively a synthetic dataset designed for 6D object pose/tracking tasks. It comprises 3D printable assembly parts and contains five evaluation sequences, each presenting different levels of difficulty for object tracking (normal (N), dynamic (D), hand occlusion (H), and blur (B)). Utilizing the GBOT dataset, we conduct a comprehensive comparison of our approach against state-of-the-art methods in 6D pose estimation performance, particularly focusing on assemblies with dynamically changing states.
@@ 3.2 Metrics
We evaluate our approach based on two main aspects: 6D pose estimation and state detection. For pose prediction, we report the absolute translation () [14], rotation error [14] and average distance error (ADD)(S) [12, 21]. ADD(S) describes the average distance error for asymmetric (ADD) and symmetric (ADD(S)).
For state detection accuracy, we calculate the F1 score. The F1 score is composed of the calculation of precision and recall for each state () and each assembly part ().
Since our goal is to utilize our state and pose recognition, we also evaluate the runtime. For the runtime, we calculate the average value over all assembly objects.
3.3 ASDF Results
We evaluate performance trade-off using different networks sizes, pose accuracy considering various refinement steps and benchmark our final approach.
3.3.1 Performance Trade-off
Since ASDF leverages object pose information to refine the state, our initial focus lies on evaluating the pose performance. As depicted in Table 3, it is evident that the fastest network performance can be attained using the smallest network configuration, which is a rational expectation. As overall network architecture we build upon YOLOv8Pose [21]. However, the underlying image backbone can vary in terms of network size. As shown in Table 3, the individual size n, m, l and xl-p6 have all their individual advantages and disadvantages for the pose performance. m and l exhibit similar performance, whereas the xl-p6 architecture, serving as our underlying backbone, demonstrates the most favorable results in terms of rotation and translation errors.
| Method | Backbone | Runtime [ms] | [mm] | erot [°] |
| No refinement | ||||
| YOLOv8Pose + state | n | 24.83 | 29.13 | 18.76 |
| YOLOv8Pose + state | xl-p6 | 50.83 | 19.11 | 11.14 |
| ICP refinment | ||||
| YOLOv8Pose + state + ICP | n | 151.77 | 12.27 | 18.96 |
| YOLOv8Pose + state +ICP | xl-p6 | 145.76 | 9.24 | 14.35 |
| Segmentation refinement | ||||
| ASDF + Seg | n | 64.46 | 8.52 | 18.76 |
| ASDF + Seg | xl-p6 | 85.03 | 6.53 | 11.14 |
| Ours | ||||
| ASDF | n | 32.45 | 8.13 | 18.76 |
| ASDF | xl-p6 | 55.70 | 5.98 | 11.14 |
| Backbone | Runtime [ms] | [mm] | erot [°] |
| n | 24.83 | 29.13 | 18.76 |
| m | 33.72 | 27.85 | 15.21 |
| l | 35.55 | 23.69 | 13.09 |
| xl-p6 | 50.83 | 19.11 | 11.14 |
@@ 3.3.1 Performance Trade-off
| Asset | ||||||||||||||||||||||
| [mm] | [mm] | [mm] | [mm] | [mm] | [mm] | [mm] | ||||||||||||||||
| N | 100.0 | 18 | 3.0 | 91.5 | 93 | 3.8 | 89.8 | 47 | 25.8 | 100.0 | 1 | 38.7 | 100.0 | 13 | 38.7 | 100.0 | 6 | 2.1 | 99.5 | 7 | 2.7 | |
| Corner | D | 100.0 | 18 | 2.7 | 99.0 | 25 | 4.8 | 88.4 | 37 | 25.0 | 100.0 | 15 | 46.8 | 100.0 | 20 | 47.8 | 100.0 | 6 | 23.2 | 100.0 | 5 | 3.5 |
| Clamp | H | 68.6 | 83 | 38.4 | 45.4 | 541 | 97.1 | 66.6 | 88 | 37.1 | 68.4 | 74 | 59.4 | 68.4 | 75 | 58.3 | 81.9 | 57 | 90.4 | 90.6 | 44 | 84.7 |
| B | 100.0 | 19 | 3.4 | 97.3 | 40 | 4.3 | 88.9 | 50 | 30.4 | 100.0 | 5 | 2.2 | 100.0 | 10 | 3.0 | 100.0 | 6 | 2.1 | 99.9 | 6 | 2.1 | |
| N | 100.0 | 4 | 4.7 | 99.6 | 14 | 10.1 | 90.6 | 18 | 5.4 | 100.0 | 2 | 2.6 | 100.0 | 3 | 3.0 | 100.0 | 2 | 2.1 | 100.0 | 5 | 3.3 | |
| Geared | D | 100.0 | 3 | 4.1 | 99.9 | 13 | 9.3 | 92.7 | 14 | 9.9 | 100.0 | 2 | 2.5 | 100.0 | 0.4 | 3.6 | 100.0 | 0.2 | 2.3 | 100.0 | 0.5 | 3.6 |
| Caliper | H | 100.0 | 8 | 6.9 | 99.2 | 24 | 12.5 | 96.5 | 9 | 4.4 | 85.4 | 30 | 30.0 | 85.5 | 31 | 30.0 | 85.4 | 30 | 30.0 | 99.6 | 8 | 7.0 |
| B | 100.0 | 4 | 4.9 | 99.6 | 14 | 9.9 | 98.9 | 9 | 8.7 | 100.0 | 2 | 2.5 | 100.0 | 3 | 2.8 | 100.0 | 2 | 2.2 | 100.0 | 5 | 3.6 | |
| N | 100.0 | 14 | 4.8 | 71.1 | 88 | 4.8 | 74.1 | 66 | 13.8 | 89.8 | 21 | 16.7 | 89.4 | 23 | 16.5 | 99.8 | 6 | 7.2 | 93.8 | 24 | 3.6 | |
| Nano | D | 90.0 | 16 | 5.0 | 72.6 | 76 | 4.6 | 63.4 | 103 | 15.6 | 87.3 | 39 | 15.3 | 87.3 | 41 | 15.8 | 96.0 | 24 | 20.0 | 92.9 | 25 | 3.7 |
| Vise | H | 98.6 | 19 | 5.3 | 65.9 | 110 | 4.7 | 61.4 | 136 | 15.3 | 76.5 | 60 | 18.3 | 75.8 | 61 | 18.1 | 72.9 | 87 | 14.5 | 87.8 | 31 | 7.3 |
| B | 100.0 | 15 | 5.4 | 70.9 | 99 | 5.2 | 61.9 | 116 | 15.2 | 91.6 | 19 | 11.4 | 91.5 | 21 | 11.1 | 95.7 | 7 | 25.1 | 92.7 | 30 | 4.7 | |
| N | 93.5 | 11 | 8.6 | 73.5 | 77 | 17.6 | 86.5 | 47 | 8.9 | 96.0 | 12 | 1.1 | 95.9 | 14 | 2.3 | 98.8 | 4 | 0.9 | 83.7 | 30 | 4.7 | |
| Screw | D | 94.5 | 11 | 9.4 | 82.0 | 67 | 18.4 | 86.4 | 56 | 27.0 | 95.9 | 17 | 2.1 | 95.9 | 22 | 3.4 | 98.8 | 6 | 1.7 | 91.6 | 21 | 6.1 |
| Clamp | H | 92.1 | 18 | 13.5 | 67.2 | 142 | 39.8 | 60.1 | 143 | 37.7 | 73.4 | 69 | 53.4 | 73.1 | 71 | 53.4 | 68.7 | 79 | 61.9 | 83.9 | 49 | 27.2 |
| B | 96.1 | 9 | 8.7 | 85.6 | 65 | 18.6 | 86.5 | 56 | 34.6 | 95.7 | 30 | 6.6 | 95.7 | 32 | 7.7 | 98.6 | 12 | 1.1 | 91.1 | 30 | 4.8 | |
| Mean | 95.8 | 17 | 8.1 | 82.52 | 93 | 16.59 | 80.8 | 62 | 19.7 | 91.3 | 25 | 19.4 | 91.2 | 28 | 19.7 | 93.5 | 21 | 17.9 | 94.2 | 20 | 10.8 | |
@@ 3.3.2 Pose Refinement
The backbone n shows the best runtime and the network size xl-p6 the best performance, see Table 3. Therefore, we further compared these two backbones considering different pose refinement methods. ASDF builds upon the use of keypoints for the 6D pose estimation. However, we considered other refinement methods as well to lift the pose performance. In Table 2 we compare the pure YOLOv8pose combined with the state prediction with the commonly used ICP-based refinement (YOLOv8pose + state + ICP). Additionally, since semantic segmentation provides clearer object boundaries, we compared our keypoint-based approach using a refinement step with a second YOLOv8 [16] semantic segmentation network and our ASDF with the backbones n and xl-p6.
3.3.3 Assembly State Detection and 6D Pose Estimation
As shown in Table 3 and Table 2, ASDF shows promising results in terms of 6D pose estimation. Comparing the assembly state detection using the F1 score, see Table 4, shows that ASDF can predict the assembly state better (79.52) compared to a pure deep learning-based approach (73.85). Moreover, as denoted in Table 4, the 6D pose performance shows an improved performance by outperforming the deep learning-based approach in ADD(S) (95.74 (ours) compare to 84.92). Moreover, for the translation error we can reduce the average error by 14.2 mm.
@@ 3.4 GBOT Results
Existing approaches face the limitation of being either used for state detection combined with 2D object detection [32] or are limited to non accessible datasets [25, 38]. However, in terms of 6D pose estimation and 6D object tracking some approaches can handle 6D pose estimation for assembled parts. We focus on multi-state assembly comparisons, therefore, we omit comparing with single state approaches such as Mb-ICG [35], which require a re-initialization per assembly step. Our comparisons builds upon the benchmark proposed by Li et al. [21].
Given that we utilize the same objects as Li et al. [21], we assess our trained models from the ASDF dataset on the GBOT dataset without the need for retraining on their training set. We exclude the comparison with LiftPod since the screws and small connectors are excluded from their evaluation due to a category-level problem. As illustrated in Table 5, our approach demonstrates superior performance compared to GBOT and their fusion of GBOT with deep learning-based re-initialization of object tracking. Moreover, in comparison to state-of-the-art tracking methods, Li et al. [21] introduce a deep learning-based approach for 6D pose estimation, which shares the same base architecture as our method. The outcomes presented in Table 5 shows that our approach, outperforms their deep learning baseline.